Required information

[The following information applies to the questions displayed below.]

Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments--Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March):

Molding Fabrication Total
Estimated total machine-hours used 2,500 1,500 4,000
Estimated total fixed manufacturing overhead $ 14,250 $ 17,550 $ 31,800
Estimated variable manufacturing overhead per machine-hour $ 3.10 $ 3.90
Job P Job Q
Direct materials $ 30,000 $ 16,500
Direct labor cost $ 34,600 $ 14,300
Actual machine-hours used:
Molding 3,400 2,500
Fabrication 2,300 2,600
Total 5,700 5,100

Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month.

Required:

For questions 1-8, assume that Sweeten Company uses a plantwide predetermined overhead rate with machine-hours as the allocation base. For questions 9-15, assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments.

4. If Job P included 20 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)
Answer #1

Calculate plantwide overhead rate :

Estimated total overhead = (2500+3.10+1500*3.90+31800) = $45400

Plantwide overhead rate = Estimated total overhead/Estimated machine hour

= 45400/4000

Plantwide overhead rate = $11.35 per machine hour

Calculate unit product Cost of job P

Job P
Direct material 30000
Direct labor 34600
Overhead applied (5700*11.35) 64695
Total product Cost 129295
Unit 20
Unit product Cost (129295/20) 6464.75